Is it possible for the equation 2x+5=0 to have a solution in the set of positive real numbers? Explain your answer.
Answer by stanbon(75887)

Is it possible for the equation 2x+5=0 to have a solution in the set of positive real numbers? Explain your answer.
--------
2x+5 = 0
2x = -5
x = -5/2
-----
Note: -5/2 is a rational number
It is therefore a Real Number.
=============
Cheers,
Stan H.